OpenStudy (anonymous):

x/7 - y/4=1 what is the x and y intercept?

OpenStudy (anonymous):

x intercept =7 y intercept is= -4

OpenStudy (debbieg):

x-intercept occurs at y=0 (on the x-axis, every y coordinate = 0) y-intercept occurs at x=0 (on the x-axis, every x coordinate = 0) So plug in x=0 and find y; plug in y=0 and find x.

OpenStudy (debbieg):

@hantenks please read the CoC: OpenStudy values the Learning process - not the ‘Give you an answer’ process. Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok so i understand that what about the =1 on the question because i have to graph it.

OpenStudy (debbieg):

To graph it, once you have the 2 intercepts, you have 2 points. So just use those for the graph.

OpenStudy (anonymous):

simply when eqn is in the form x/a +y/b =1 a, b are x,y intercepts respectively i.e points are (0,b) and (a.0)

OpenStudy (anonymous):

join them

OpenStudy (debbieg):

To find the intercepts, like I said: The x-intercept is where the graph crosses the x-axis, so the y-coordinate = 0 x/7 - 0/4=1 x/7=1 Solve for x, by multiplying both sides by 7 That is the x-coordinate of the x-intercept, and the y-coordinate for that point is =0 similarly for the y-intercept, but with x=0.

OpenStudy (anonymous):

ok i get what your saying thank you

OpenStudy (anonymous):

@ debbiG whats Coc??

OpenStudy (debbieg):

The site Code of Conduct. I linked it above.

OpenStudy (anonymous):

@DebbieG thanx..i am new to this